Jerome blends Brand A of coffee that sells for $9.50 a pound with Brand B of coffee that sells for $12.00 a pound. If the 45 pounds of blend sells for $10.50 a pound, how many pounds of each type of coffee did Jerome use?
Answer by checkley77(12844)

12x+9.50(45-x)=10.50*45
12x+427.5-9.5x=472.5
2.5x=472.5-427.5
2.5x=45
x=45/2.5
x=18 pounds of the $12 coffee is used.
45-18=27 pounds of the $9.50 coffee is used.
Proof:
12*18+9.5*27=10.50*45
216+256.5=472.5
472.5=472.5